# DAmodel: Hierarchical Bayesian Modelling of DA White Dwarfs for Spectrophotometric Calibration

### v1

@@ 5 Future Work
Aside from the purpose of calibration, our novel statistical framework opens doors to population analysis on larger samples of DA WDs. There are nearly 2000 DESI DA WDs (Manser et al., 2024) and a further 500 DA WDs in Gaia (O’Brien et al., 2024), with both sets having spectroscopic and photometric coverage. A volume limited subset of these samples could be analysed by our framework to provide more meaningful statistics on population dust parameters. Alternatively, we could take the same approach as Fitzpatrick & Massa (1990) by adding additional parameters to the extinction function to infer our own dust relation. We could also extend the framework to provide population inferences on the surface gravity and effective temperature parameters. If we introduce Gaia parallaxes to our spectrophotometric framework, we could use the inferred surface gravity to estimate mass evolution of DA white dwarfs (Si et al., 2017, 2018). These ideas would not have been computationally feasible in the past, but are now within reach thanks to our GPU accelerated implementation of DAmodel.
An important science case that motivates the need for our all-sky accurate flux standards is SN Ia cosmology. Most recent efforts to constrain our Universe use compilations of SNe from a variety of surveys (Scolnic et al., 2022; Brout et al., 2022; Rubin et al., 2023). An essential part of the analysis is the cross-calibration of the surveys so they are on a consistent flux scale (Scolnic et al., 2015). Popovic et al. (2025) are already using our inferred SEDs in their Dovekie cross-calibration codebase. Their work will combine the present day’s largest SNe Ia samples from the Zwicky Transient Facility (Rigault et al., 2024), Dark Energy Survey (Abbott et al., 2024) and more, to further constrain the nature of our Universe.
@@ Acknowledgements
BMB is supported by the Cambridge Centre for Doctoral Training in Data-Intensive Science funded by the UK Science and Technology Facilities Council (STFC) and two G-Research early career researchers grants used for equipment and travel. BMB also acknowledges travel support provided by Christ’s College Cambridge’s Clayton Fund, LSST DESC supported by DOE funds administered by SLAC and STFC for UK participation in LSST through grant ST/S006206/1. BMB, AB, ML and GN were supported on this project through the HST Guest Observer Program (HST-GO-16764, PI G. Narayan). GN also gratefully acknowledges NSF support from AST-2206195, and a CAREER grant AST-2239364, supported in-part by funding from Charles Simonyi, and OAC-2311355, DOE support through the Department of Physics at the University of Illinois, Urbana-Champaign (#13771275), and HST-GO-17128 (PI R. Foley). Supernova and astrostatistics research at Cambridge University is supported by the European Union’s Horizon 2020 research and innovation programme under European Research Council Grant Agreement No 101002652 (PI K. Mandel) and Marie Skłodowska-Curie Grant Agreement No 873089.
@@
The implementation of DAmodel is publically available at: https://github.com/benboyd97/DAmodel. All data used and our published SEDs can be found at: https://zenodo.org/records/14339961. The theoretical DA WD spectra (Hubeny et al., 2021) used to create the underlying grid can be found at: www.as.arizona.edu/~hubeny/pub/DAgrid24.tar.gz. The SEDs used to tie our work to the CALSPEC system (Bohlin et al., 2020) may be found at: https://archive.stsci.edu/hlsps/reference-atlases/cdbs/current_calspec/. The HST/WFC3 UVIS2 and IR filters (Calamida et al., 2022a) used in this work are available at: https://www.stsci.edu/hst/instrumentation/wfc3/performance/throughputs. Data from other surveys used to calculate residuals is published in Axelrod et al. (2023).
@@ Appendix A Published SEDs
In Figure 12 we illustrate the SEDs for each WD in our network that we publish alongside our work. Each published WD SED is derived by averaging 2000 randomly drawn SEDs forward-modelled from 2000 posterior samples. With these SED draws we may also find the associated uncertainty at each wavelength found by taking the standard deviation at each pixel. Our published SEDs can be found at: https://zenodo.org/records/14339961.
@@
Axelrod et al. (2023) previously analysed the same network of stars. We plot the average ratio between our inferred SEDs and those from Axelrod et al. (2023) at wavelengths with mutual coverage in Figure 13. The largest differences are observed in the UV, likely driven by our inference of the $R_{V}$ parameter, the adoption of the updated dust relation and theoretical template, and the inclusion of HST/STIS UV spectra. Spikes in the optical are due to differences in inferred effective temperature, resulting in differing absorption feature strength.
@@ Appendix E Population Dust Constraints
As a demonstration of our models ability to infer population-level parameters, we infer a $\mu_{RV}$ and $\sigma_{RV}$ given a truncated normal prior $R^{s}_{V}\sim TN(\mu_{RV},\sigma_{RV},1.2,\infty)$. We also infer an average population extinction $\tau$ assuming $A^{s}_{V}$ follows an exponential distribution. The resulting posteriors with and without using the HST/STIS UV spectra are shown in Figure 18. The inferred parameters are significantly different when compared to the constant population parameters used in our main analysis motivated by the work of Schlafly et al. (2016). The results in our population inference are to be taken with caution, as our DA WD network is strongly influenced by selection effects that are not taken into account in the analysis. In Section 5 we discuss future plans to apply the model to a larger volume limited sample of DA WDs to provide constraints on these population parameters that are uninfluenced by selection effects. The results show that the introduction of the HST/STIS UV spectra does not significantly change the $R_{V}$ population parameters. We would only expect substantial difference if we obtain NIR spectra for a panchromatic view to monitor the effects of the dust relations.
@@ Appendix H Residuals with different modelling assumptions
Table 5: HST/WFC3 residual mean $\mu$ and RMS for the 32 faint DA WDs when comparing their observed photometry with synthetic photometry for different experiments. Opt only signifies that we use only ground-based optical spectra and HST photometry to calibrate the stars. Opt+STIS signifies that we also use the STIS UV spectra for the 19 WDs that had coverage. G23 indicates that we use the Gordon et al. (2023) dust relation, whilst F99 represents the Fitzpatrick (1999) dust relation. Const $R_{V}$ refers to a constant $R_{V}=3.1$. Norm $R_{V}$ experiments used the truncated normal $R_{V}$ prior $TN(\mu_{RV}=3.1,\sigma_{RV}=0.18,1.2,\infty)$. Population $R_{V}$ experiments also inferred the population parameters $\mu_{RV}$, $\sigma_{RV}$ and $\tau$. We label tests presented in our final results with an *. Results shown are in AB millimagnitudes.
@@ Appendix I Photometric Residuals with different surveys
Using our inferred synthetic SEDs for the WDs, we can integrate over various bandpass filters corresponding to different surveys. For any survey that observed a given standard, we can compute the residual between their measurement and our synthetic prediction. We show our residuals for DES in Figure 9, SDSS in Figure 21, PS1 in Figure 22, Gaia in Figure 23 and DECaLS in Figure 24. The data collection process and residual results are discussed in Section 4.4. Since each survey is calibrated independently, a wavelength-dependent bias is expected. Additional biases are anticipated because our SEDs are tied to the new CALSPEC Bohlin et al. (2020) system rather than Bohlin (2014). Quantifying these biases is crucial for cross-calibrating surveys and combining samples. In Section 5, we discuss the ongoing application of our SEDs to unify SN Ia samples across different surveys.
Figure 21: Photometric residuals when comparing our synthetic SDSS photometry with real SDSS observations (Holberg & Bergeron, 2006) for the faint DA WD standards that had coverage. Objects are arranged in order of brightness, ending with the dimmest object. We presented the unweighted mean and RMS residual for each band. Residuals are in AB magnitudes.
Figure 22: Photometric residuals when comparing our synthetic PS1 photometry with real PS1 observations (Flewelling et al., 2020) for the faint DA WD standards that had coverage. Objects are arranged in order of brightness, ending with the dimmest object. We presented the unweighted mean and RMS residual for each band. Residuals are in AB magnitudes.
Figure 23: Photometric residuals when comparing our synthetic Gaia photometry with real Gaia observations (Gaia Collaboration et al., 2023) for the faint DA WD standards that had coverage. Objects are arranged in order of brightness, ending with the dimmest object. We presented the unweighted mean and RMS residual for each band. Residuals are in Vega magnitudes.
Figure 24: Photometric residuals when comparing our synthetic DECaLS photometry with real DECaLS observations (Schlegel et al., 2021) for the faint DA WD standards that had coverage. Objects are arranged in order of brightness, ending with the dimmest object. We presented the unweighted mean and RMS residual for each band. Residuals are in AB magnitudes.

### v2

@@ 5 Future Work
Aside from the purpose of calibration, our novel statistical framework opens doors to population analysis on larger samples of DA WDs. There are nearly 2000 DESI DA WDs (Manser et al., 2024) and a further 500 DA WDs in the Gaia 40 pc sample (O’Brien et al., 2024), with both sets having spectroscopic and photometric coverage. Analysing these volume-limited subsets with our framework could provide more meaningful statistics on population dust parameters. Alternatively, we could take the same approach as Fitzpatrick & Massa (1990) by adding additional parameters to the extinction function to infer our own dust relation. We could also extend the framework to provide population inferences on the surface gravity and effective temperature parameters. If we introduce Gaia parallaxes to our spectrophotometric framework, we could use the inferred surface gravity to estimate mass evolution of DA white dwarfs (Si et al., 2017, 2018). These ideas would not have been computationally feasible in the past, but are now within reach thanks to our GPU accelerated implementation of DAmodel.
An important science case that motivates the need for our all-sky accurate flux standards is SN Ia cosmology. Most recent efforts to constrain our Universe use compilations of SNe from a variety of surveys (Scolnic et al., 2022; Brout et al., 2022; Rubin et al., 2023). An essential part of the analysis is the cross-calibration of the surveys so they are on a consistent flux scale (Scolnic et al., 2015). Popovic et al. (2025, in prep.) are already using our inferred SEDs in their Dovekie cross-calibration codebase. Their work will combine the present day’s largest SNe Ia samples from the Zwicky Transient Facility (Rigault et al., 2025), Dark Energy Survey (des2024) and more, to further constrain the nature of our Universe.
@@ Acknowledgements
BMB is supported by the Cambridge Centre for Doctoral Training in Data-Intensive Science funded by the UK Science and Technology Facilities Council (STFC) and two G-Research early career researchers grants used for equipment and travel. BMB also acknowledges travel support provided by Christ’s College Cambridge’s Clayton Fund, LSST DESC supported by DOE funds administered by SLAC and STFC for UK participation in LSST through grant ST/S006206/1. BMB, AB, ML and GN were supported on this project through the HST Guest Observer Program (HST-GO-16764, PI G. Narayan). GN also gratefully acknowledges NSF support from AST-2206195, and a CAREER grant AST-2239364, supported in-part by funding from Charles Simonyi, and OAC-2311355, DOE support through the Department of Physics at the University of Illinois, Urbana-Champaign (#13771275), and HST-GO-17128 (PI R. Foley). Supernova and astrostatistics research at Cambridge University is supported by the European Union’s Horizon 2020 research and innovation programme under European Research Council Grant Agreement No 101002652 (PI K. Mandel) and Marie Skłodowska-Curie Grant Agreement No 873089. CWS acknowledges support from Harvard University and the DOE cosmic frontier program through award DE-SC0007881. EO was also partially supported by NSF grants AST-1815767 and AST-1313006. We acknowledge support from STSCI/HST: HST-GO-12967, HST-GO-13711, HST-GO-15113.
Spectroscopic observations reported here were obtained at the MMT Observatory, a joint facility of the University of Arizona and the Smithsonian Institution. Spectra were also obtained at SOAR and at Gemini based in part on observations obtained at the Southern Astrophysical Research (SOAR) Telescope, which is a joint project of the Ministério da Ciência, Tecnologia e Inovações do Brasil (MCTI/LNA), the US National Science Foundation’s NOIRLab, the University of North Carolina at Chapel Hill (UNC), and Michigan State University (MSU). Observations were also obtained at the international Gemini Observatory, a program of NSF’s NOIRLab, which is managed by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the National Science Foundation on behalf of the Gemini Observatory partnership: the National Science Foundation (United States), National Research Council (Canada), Agencia Nacional de Investigación y Desarrollo (Chile), Ministerio de Ciencia, Tecnología e Innovación (Argentina), Ministério da Ciência, Tecnologia, Inovações e Comunicações (Brazil), and Korea Astronomy and Space Science Institute (Republic of Korea). This work makes use of observations from the Las Cumbres Observatory global telescope network. This research has made use of the SVO Filter Profile Service "Carlos Rodrigo", funded by MCIN/AEI/10.13039/501100011033/ through grant PID2023-146210NB-I00.
AS dedicates his efforts towards this project to the memory of his late advisor Dr. J.B. Oke, whose measurements of the absolute flux distribution of Vega pioneered the application of spectrophotometry in astronomy.
@@
The implementation of DAmodel v1.0 is publically available at: https://github.com/benboyd97/DAmodel. All data used, coordinates and our published v1.0 SEDs can be found at: https://zenodo.org/records/14339960. The theoretical DA WD spectra (Hubeny et al., 2021) used to create the underlying grid can be found at: www.as.arizona.edu/~hubeny/pub/DAgrid24.tar.gz. The SEDs used to tie our work to the CALSPEC system (Bohlin et al., 2020) may be found at: https://archive.stsci.edu/hlsps/reference-atlases/cdbs/current_calspec/. The HST/WFC3 UVIS2 and IR filters (Calamida et al., 2022a) used in this work are available at: https://www.stsci.edu/hst/instrumentation/wfc3/performance/throughputs. Data from other surveys used to calculate residuals is published in Axelrod et al. (2023).
@@ Appendix A Published SEDs
In Figure 12 we illustrate the SEDs for each WD in our network that we publish alongside our work. Each published WD SED is derived by averaging 2000 randomly drawn SEDs forward-modelled from 2000 posterior samples. With these SED draws we may also find the associated uncertainty at each wavelength found by taking the standard deviation at each pixel, shown in Figure 13. Our published SEDs and associated errors, as well as coordinates, can be found at: https://zenodo.org/records/14339960.
@@
Figure 13: Inferred SED flux error fraction as a function of wavelength for the DA WDs in our network. These errors are estimated by sampling the inferred WD parameters to generate random realisations of the SEDs, with the standard deviation at each pixel representing the uncertainty. We use the average of the sampled SEDs to determine the flux at each pixel.
Table 4: Resources released with this work, along with those utilised in its methodology.
Resource Location
Full coordinates of the 35 DA WDs https://zenodo.org/records/14939532/files/wdfs_coords.txt
35 DA WD SEDs and corresponding errors derived from v1.0 of this work, along with all spectroscopic and photometric data utilised https://zenodo.org/records/14339960
Implementation of DAmodel v1.0 used in this work https://github.com/benboyd97/DAmodel
Theoretical DA WD spectra computed with Tlusty v208 (Hubeny et al., 2021) www.as.arizona.edu/~hubeny/pub/DAgrid24.tar.gz
SEDs used to tie the network to the CALSPEC system: g191b2b_mod_012.fits, gd153_mod_012.fits and gd71_mod_012.fits (Bohlin et al., 2020) https://archive.stsci.edu/hlsps/reference-atlases/cdbs/current_calspec/
HST/WFC3 UVIS2 and IR filters (Calamida et al., 2022a) https://www.stsci.edu/hst/instrumentation/wfc3/performance/throughputs
Data from other surveys used to calculate residuals Published in Axelrod et al. (2023)
@@
Axelrod et al. (2023) previously analysed the same network of stars. We plot the average ratio between our inferred SEDs and those from Axelrod et al. (2023) at wavelengths with mutual coverage in Figure 14. The largest differences are observed in the UV, likely driven by our inference of the $R_{V}$ parameter, the adoption of the updated dust relation and theoretical template, and the inclusion of HST/STIS UV spectra. Spikes in the optical are due to differences in inferred effective temperature, resulting in differing absorption feature strength. In Figure 16 we show the difference in synthetic HST photometry between our work and Axelrod et al. (2023). The residuals are mostly driven by the adoption of the updated Bohlin et al. (2020) CALSPEC system, whilst mmag-level deviations are caused by differences in modelling choices and methodology.
@@
Figure 16: Residuals of synthetic HST photometry between our work and the previous work of Axelrod et al. (2023) for the full network of 35 WDs. The work of Axelrod et al. (2023) was tied to the older Bohlin (2014) system. The coloured residuals and statistics are using our published synthetic photometry that is tied the updated Bohlin et al. (2020) CALSPEC system. In white are the residuals when we use the Bohlin (2014) CALSPEC system that matches what was used in Axelrod et al. (2023). Residuals shown are in AB mags and error bars represent the standard deviation of the inferred synthetic magnitudes from our work only. The unweighted mean and RMS scatter do not include the three primary standards that are illustrated by the star markers.
@@ Appendix E Population Dust Constraints
As a demonstration of our models ability to infer population-level parameters, we infer a $\mu_{RV}$ and $\sigma_{RV}$ given a truncated normal prior $R^{s}_{V}\sim\mathcal{TN}(\mu_{RV},\sigma_{RV},1.2,\infty)$. We also infer an average population extinction $\tau$ assuming $A^{s}_{V}$ follows an exponential distribution. The resulting posteriors with and without using the HST/STIS UV spectra are shown in Figure 20. The inferred parameters are significantly different when compared to the constant population parameters used in our main analysis motivated by the work of Schlafly et al. (2016). The results in our population inference are to be taken with caution, as our DA WD network is strongly influenced by selection effects that are not taken into account in the analysis. In Section 5 we discuss future plans to apply the model to a larger volume limited sample of DA WDs to provide constraints on these population parameters that are uninfluenced by selection effects. The results show that the introduction of the HST/STIS UV spectra does not significantly change the $R_{V}$ population parameters. We would only expect substantial difference if we obtain NIR spectra for a panchromatic view to monitor the effects of the dust relations.
@@ Appendix H Residuals with different modelling assumptions
Table 6: HST/WFC3 residual mean $\mu$ and RMS for the 32 faint DA WDs when comparing their observed photometry with synthetic photometry for different experiments. Opt only signifies that we use only ground-based optical spectra and HST photometry to calibrate the stars. Opt+STIS signifies that we also use the STIS UV spectra for the 19 WDs that had coverage. G23 indicates that we use the Gordon et al. (2023) dust relation, whilst F99 represents the Fitzpatrick (1999) dust relation. Const $R_{V}$ refers to a constant $R_{V}=3.1$. Norm $R_{V}$ experiments used the truncated normal $R_{V}$ prior $\mathcal{TN}(\mu_{RV}=3.1,\sigma_{RV}=0.18,1.2,\infty)$. Population $R_{V}$ experiments also inferred the population parameters $\mu_{RV}$, $\sigma_{RV}$ and $\tau$. We label tests presented in our final results with an *. Results shown are in AB millimagnitudes.
@@ Appendix I Photometric Residuals with different surveys
Using our inferred synthetic SEDs for the WDs, we can integrate over various bandpass filters corresponding to different surveys. For any survey that observed a given standard, we can compute the residual between their measurement and our synthetic prediction. We show our residuals for DES in Figure 9, SDSS in Figure 23, PS1 in Figure 24, Gaia in Figure 25 and DECaLS in Figure 26. The data collection process and residual results are discussed in Section 4.4. Since each survey is calibrated independently, a wavelength-dependent bias is expected. Additional biases are anticipated because our SEDs are tied to the new CALSPEC Bohlin et al. (2020) system rather than Bohlin (2014). In the Figures we show that the residuals are significantly smaller, comparable to those in Axelrod et al. (2023), when we tie our network to the the older CALSPEC system (Bohlin, 2014). In Figure 25 we also find smaller residuals when using the same scaled definition of Vega as Gaia (Riello et al., 2021). Despite getting better agreement with other surveys when using the older CALSPEC system, we maintain our decision to publish the network with respect to the more accurate Bohlin et al. (2020) CALSPEC system and encourage other surveys to adopt the same for their absolute calibration. Quantifying these biases is crucial for cross-calibrating surveys and combining samples. In Section 5, we discuss the ongoing application of our SEDs to unify SN Ia samples across different surveys.
Figure 23: Photometric residuals when comparing our synthetic SDSS photometry with real SDSS observations (Holberg & Bergeron, 2006) for the faint DA WD standards that had coverage. Objects are arranged in order of brightness, ending with the dimmest object. We presented the unweighted mean and RMS residual for each band. Residuals are in AB magnitudes. Plotted underneath in white are the residuals when using the Bohlin (2014) CALSPEC system rather than the Bohlin et al. (2020) system.
Figure 24: Photometric residuals when comparing our synthetic PS1 photometry with real PS1 observations (Flewelling et al., 2020) for the faint DA WD standards that had coverage. Objects are arranged in order of brightness, ending with the dimmest object. We presented the unweighted mean and RMS residual for each band. Residuals are in AB magnitudes. Plotted underneath in white are the residuals when using the Bohlin (2014) CALSPEC system rather than the Bohlin et al. (2020) system.
Figure 25: Photometric residuals when comparing our synthetic Gaia photometry with real Gaia observations (Gaia Collaboration et al., 2023) for the faint DA WD standards that had coverage. Objects are arranged in order of brightness, ending with the dimmest object. We presented the unweighted mean and RMS residual for each band. Residuals are in Vega magnitudes. Plotted in white are the residuals when using the Bohlin (2014) CALSPEC system (rather than the Bohlin et al. (2020) system) and Gaia’s scaled definition of Vega.
Figure 26: Photometric residuals when comparing our synthetic DECaLS photometry with real DECaLS observations (Schlegel et al., 2021) for the faint DA WD standards that had coverage. Objects are arranged in order of brightness, ending with the dimmest object. We presented the unweighted mean and RMS residual for each band. Residuals are in AB magnitudes. Plotted in white are the residuals when using the Bohlin (2014) CALSPEC system rather than the Bohlin et al. (2020) system.